OpenStudy (anonymous):

Multiply, and then simplify the problem. Assume that all variables represent positive real numbers. [(√2+√3)-√6] [(√2+√3)+√6]

OpenStudy (kc_kennylau):

variable?

OpenStudy (kc_kennylau):

use the difference of squares identity :D (identity: \(a^2-b^2=(a+b)(a-b)\))

OpenStudy (anonymous):

I should use FOIL on the binomials, and then what am I suppose to do with the -√6 and √6?

OpenStudy (kc_kennylau):

no you do this in your mind: a=\(\sqrt2+\sqrt3\), b=\(\sqrt6\)

OpenStudy (kc_kennylau):

\([(\sqrt2+\sqrt3)-\sqrt6][(\sqrt2+\sqrt3)+\sqrt6]\) =\((\sqrt2+\sqrt3)^2-\sqrt6^2\)

OpenStudy (anonymous):

Is that how it would be simplified?

OpenStudy (kc_kennylau):

this is just the first step :D (do you get the first step)

OpenStudy (anonymous):

I see how you got that, but how does it end up being: 2√6-1 in the answer key?

OpenStudy (kc_kennylau):

why would it not? :)

OpenStudy (anonymous):

Do I use foil to get that 2? I know that √2*√2=2. I thought that was how they got that part, but I don't see where the -1 comes from.

OpenStudy (kc_kennylau):

don't let the answer key affect you

OpenStudy (kc_kennylau):

\([(\sqrt2+\sqrt3)-\sqrt6][(\sqrt2+\sqrt3)+\sqrt6]\) =\((\sqrt2+\sqrt3)^2-\sqrt6^2\) =\((\sqrt2^2+2\sqrt2\sqrt3+\sqrt3^2)-6\)

OpenStudy (kc_kennylau):

do you actually remember the identities

OpenStudy (kc_kennylau):

Let me type the idens all out

OpenStudy (kc_kennylau):

\(1.\) \((a+b)^2=a^2+2ab+b^2\) \(2.\) \((a-b)^2=a^2-2ab+b^2\) \(3.\) \(a^2-b^2=(a+b)(a-b)\) \(4.\) \(a^3+b^3=(a+b)(a^2-ab+b^2)\) \(5.\) \(a^3-b^3=(a-b)(a^2+ab+b^2)\)

OpenStudy (kc_kennylau):

so now we have: \([(\sqrt2+\sqrt3)-\sqrt6][(\sqrt2+\sqrt3)+\sqrt6]\) =\((\sqrt2+\sqrt3)^2-\sqrt6^2\) =\((\sqrt2^2+2\sqrt2\sqrt3+\sqrt3^2)-6\) could you continue? :)

OpenStudy (anonymous):

I wouldn't know what to do with that. Am I suppose to somehow simplify that further? Or do I just do something with the -6 first?

OpenStudy (kc_kennylau):

simplify that further :)

OpenStudy (anonymous):

How do I simplify it further? Is it: (2√2^3+√3^3)-6?

OpenStudy (kc_kennylau):

nope

OpenStudy (kc_kennylau):

what is the definition of sqrt?

OpenStudy (kc_kennylau):

ok let's do it terms by term

OpenStudy (kc_kennylau):

look at what we got here: \((\sqrt2^2+2\sqrt2\sqrt3+\sqrt3^2)-6\) The first term \(\sqrt2^2\) is equivalent to?

OpenStudy (anonymous):

2

OpenStudy (kc_kennylau):

and the second term \(2\sqrt2\sqrt3\) is?

OpenStudy (anonymous):

2√6

OpenStudy (kc_kennylau):

and the third?

OpenStudy (anonymous):

3

OpenStudy (kc_kennylau):

and the whole thing?

OpenStudy (anonymous):

I see it now! It's: 5+2√6-6, which turns into: 2√6-1. That's great, thank you!

OpenStudy (kc_kennylau):

no problem :)